b Quark Tagging in W+Jet Events

Applied these tagging techniques to W+jet events

Sample  SVX bkg SVX tags  SLT bkg  SLT tags
W+≥ 3 jet  6.7 ± 2.1  27  15.4 ± 2.3  23

• Probability of background fluctuation is 1 × 10−6 – Combining SVX, SLT and DIL channels

• Agrees well with expected rate from t t production

SVX Tagging Results

The SVX tagged sample is single most significant observation

• Observe 27 tags when we expect 6.7 ± 2.1 background • Probability of background fluctuation 2 × 10−5